OpenStudy (anonymous):

Consider a standard uniform density. The mean for this density is .5 and the variance is 1 / 12. You sample 1,000 sample means where each sample mean is comprised of 100 observations. You take the standard deviation of the 1,000 sample means. About what number would you expect it to be? Give your answer to at least 3 decimal places.

OpenStudy (kropot72):

The standard deviation of the sample mean is given by \[\frac{\sigma}{\sqrt{n}}=\frac{\sqrt{\frac{1}{12}}}{\sqrt{100}}\] where sigma is the standard deviation of the population and n is the sample size.

OpenStudy (anonymous):

THX

OpenStudy (kropot72):

You're welcome :)